# Unusual low-temperature behavior in the half-filled band of the one-dimensional extended Hubbard model in atomic limit

### v1

@@ I Introduction
Recent studies on various effective-one-dimensional models with short-range interactions have revealed intriguing thermal behaviors, resembling first and second-order phase transitions [1]. This peculiar behavior, especially around pseudo-transitions, has been further analyzed in reference [2], focusing on spin correlation functions. The simplest models where this unusual behavior arises are in the Ising diamond chain [3, 4, 5]. Another similar decorated Ising chain discussed is the Ising sawtooth-like chain model [5], along with two and three-leg Ising ladder models [6, 7, 8]. Pseudo-transitions also appear in other models, such as the Ising-Heisenberg diamond chain [9, 10], and the one-dimensional double-tetrahedral model with alternating Ising spins and delocalized electrons [11]. Similar phenomena are observed in ladder models with Ising-Heisenberg coupling [12] and triangular tube models [13], highlighting a pattern of pseudo-transition. Further investigations [6, 14, 15] have shown some kind of universality of pseudo-critical exponents, while still satisfying the Rushbrooke inequality. All the above models are related to classical and classical-quantum spin models; here by classical we mean that the Hamiltonian does not contain non-commuting terms. However, models of other natures, like the extended Hubbard model in a diamond chain structure [16], and the Potts model on a diamond chain structure, have also exhibited this unusual behavior. On the other hand, exploring pseudo-transitions in genuine one-dimensional systems without decoration couplings is rather interesting. In this sense, the Zimm-Bragg-Potts model was recently explored and found to exhibit this anomalous behavior [17].
@@
On the other hand, the Hubbard model [18] stands as a foundational model in the modern theory of strongly correlated electrons, playing a paradigmatic role in the study of electronic correlations in quantum materials. This model is particularly significant in contexts where interactions are crucial. The Bethe ansatz method helps to understand this model better, especially in figuring out how electrons behave in different conditions and how they interact with each other [19]. The one-dimensional extended Hubbard model is a simplified theoretical model that describes the behavior of electrons in a one-dimensional chain, which additionally includes nearest-neighbor interaction energy terms. The nearest-neighbor interaction energy term describes the Coulomb repulsion between electrons occupying neighboring sites. The one-dimensional extended Hubbard model has been widely studied in the literature due to its relevance in understanding the electronic properties of one-dimensional materials such as carbon nanotubes and organic conductors [20]. It also serves as a simple prototype model for studying strongly correlated electron systems [21]. On the other hand, several theoretical investigations have focused on the extended Hubbard model in the one-dimensional case. Numerous investigations regarding the half-filled ground states of the extended Hubbard model have been conducted. Tsuchiizu and Furusaki [22] revisited the ground-state phase diagram of the one-dimensional half-filled extended Hubbard model, revealing new phases and transitions using a renormalization-group approach. Glocke et al. [23] utilized density-matrix renormalization group methods on transfer matrices to study the thermodynamics of the one-dimensional extended Hubbard model at half-filling. This highlights the detection of phase transitions through standard thermodynamic measures like isothermal compressibility and magnetic susceptibility. The study identifies a unique phase with long-range dimer order and delineates the phase diagram, comparing it with quantum Monte Carlo studies [24, 25]. The phase diagram and critical exponents of the one-dimensional $U-V$ model at quarter-filling were identified using exact diagonalization and various limit results, identifying a transition from Luttinger liquid to charge density wave insulator, noting dominant superconducting or spin density wave fluctuations depending on the value of $V$ [26]. Further investigation was also performed in reference [27].
@@
The present work is organized as follows: In Sec. 2, we give the thermodynamics of the one-dimensional extended Hubbard model in the atomic limit, identifying each eigenvalue of the cubic root and determining which one is the largest, a topic not previously elucidated. In Sec. 3, we analyze a peculiar property in the low-temperature region and explore the so-called pseudo-critical temperature transition, along with the corresponding region where the main properties undergo significant changes at this anomalous temperature. This phenomenon is what we refer to as the quasi-phase diagram [1]. In Sec. 4, we explore additional physical quantities, reporting the influence of pseudo-transition. Finally, in Sec. 5, we present our conclusions.
@@ II.2 Thermodynamic quantities and correlators
Here, we introduce several useful thermodynamic quantities. The entropy is calculated as ${\cal S}=-\frac{\partial\Omega}{\partial T}$; the internal energy is given by $\mathcal{E}=-\frac{\partial\ln\left(\lambda_{0}\right)}{\partial\beta}=k_{B}T^{2}\frac{\partial\ln\left(\lambda_{0}\right)}{\partial T}$; analogously we can obtain specific heat at constant chemical potential $C=T\frac{\partial{\cal S}}{\partial T}=\frac{\partial{\cal E}}{\partial T}$ [38]; the magnetization can be expressed as $M=-\frac{\partial\Omega}{\partial h}$; magnetic susceptibility $\chi=\frac{\partial M}{\partial h}$; the electron density $\rho$ is determined by $\rho=-\frac{\partial\Omega}{\partial\mu}$; and the isothermal compressibility is derived from $\kappa=\tfrac{1}{\rho^{2}}\frac{\partial\rho}{\partial\mu}$.
@@ III Phase diagrams
In what follows, we discuss the ground-state energy, focusing on peculiar regions where anomalous behavior appears. Then we consider the low-temperature case to discuss quasi-phase diagrams and define a pseudo-critical temperature.
@@ III.1 Zero-temperature phase diagram
Table 1: The second column provides zero-temperature phase boundary conditions, the third column gives the chemical potential at the phase boundary, while the fourth column reports the associated residual boundary entropy.
@@ III.2 Anomalous behavior in low-temperature region and quasi-phase diagram
Figure 5: Quasi-phase diagram depicting the half-filled band in the $h/U-T/U$ plane, with fixed parameters $V/U=0.54$ and $\mu/U=1.58$. The dashed curve delineates the pseudo-critical-temperature $T_{p}$. Panels (a) and (b) showcase density plots of magnetization and entropy, respectively.
@@
In exploring another intriguing aspect of the low-temperature region, Fig. 5a presents a density plot of magnetization $M$ in the $h/U-T/U$ plane, assuming $V/U=0.54$ and $\mu/U=1.58$. Here, the dashed line represents the pseudo-critical temperature $T_{p}$. Below a magnetic field of $h/U=0.04$ at sufficiently low temperatures, the system resides in a quasi alternating-pair $qAP$ region, characterized by particles predominantly arranged in alternating pairs. As the magnetic field increases within this low temperature regime, the system goes swiftly into a quasi-paramagnetic $qPM$ region, where most spins, one per site, align parallel to the magnetic field. Conversely, at very weak magnetic fields and increasing temperatures, the $FR_{2}$ frustrated system leads into a quasi-frustrated $qFR_{2}$ region. The boundary between $qFR_{2}$ and $qPM$ is marked by a standard crossover, or as visualized in the magnetization plot, a diffuse boundary, appearing as a fuzzy region. In panel (a), one can observe the magnetization shift between $qAP$ and $qPM$ regions, although the boundary between $qAP$ and $qFR_{2}$ remains indistinct due to null magnetization in both regions. Similarly, Fig. 5b displays the entropy $\mathcal{S}$, using the same parameters as panel (a). Here, the crossover boundary, denoted by a dashed line, outlines the pseudo-critical temperature as a function of magnetic field. In contrast, panel (b) shows that the boundary between $qAP$ and $qPM$ becomes indistinguishable with almost null entropy, while the boundary between $qAP$ and $qFR_{2}$ is sharply defined. This distinction is due to the residual entropy of $FR_{2}$ being $\mathcal{S}=k_{B}\ln(2)$. However, at temperatures above $T>0.06$, the entropy significantly increases, becoming clearly larger than $\mathcal{S}\gtrsim 0.7k_{B}$.
III.3 Pseudo-critical temperature condition
@@
This detailed analysis further explores the interesting behavior in the low-temperature region of the model. Previously, the pseudo-critical temperature was defined for models where the free energy involved a simple square root expression [1, 2, 14]. Defining the pseudo-critical temperature generally poses a challenge, as the only available indicator is an anomalous behavior in the low-temperature region, marked by significant changes in entropy and magnetization, along with sharp peaks in correlation length, specific heat, and magnetic susceptibility. At very low temperatures, these peaks occur at roughly the same temperature. However, as the anomalous behavior shifts to higher temperatures $T_{p}$, these peaks appear at slightly different temperatures for each quantity, complicating the definition of the pseudo-critical temperature.
@@
An interesting observation is that this anomalous behavior is evident in fundamental quantities such as the ratio between the second largest and the largest eigenvalues. To analyze this more deeply, consider the temperature and magnetic field dependence of the eigenvalues. At a fixed magnetic field $h_{1}$, there exists a special temperature, the pseudo-critical temperature $T_{p}$, which satisfies the relation
@@
Panel (b) of the Fig. 6 illustrates the function $g(T,h)$ as it varies with temperature, under the conditions of a null magnetic field, $\mu=1.58$, $U=1$, and $V=0.55$. In this representation, one can observe that the function $g(T,h)$ passes through zero at a specific temperature, denoted as $T_{1}$. This provides a clear visual indication of the pseudo-critical point in the behavior of the system.
@@
Furthermore, by employing a perturbative approach, as elaborated in Appendix A and illustrated through Eqs. (41) and (42) for the unperturbed result, we reaffirm the critical equation derived earlier, simplifying our findings into this coherent expression.
In regions of low temperature, $T_{p}\approx T_{p^{*}}\approx T_{1}$, which are very close to each other. When this occurs, we can observe anomalous behavior, and we simply define it as the pseudo-critical temperature, $T_{p}$.
For the present model the condition (36) provides a sufficiently accurate condition for determining the pseudo-critical temperature, where $T_{p}\approx T_{1}$ holds. Specifically, the pseudo-critical temperature for a fixed magnetic field $h_{1}$ can be represented using the results $w_{0,2}={\rm e}^{\beta(\mu-U/2)}$, $w_{1,1}={\rm e}^{\beta(\mu-V)}$, and $z=2\cosh(\beta h)$, leading to the condition
@@
In Fig. 7, the temperatures $T_{1}$ (represented by a thin black line), $T_{p^{*}}$ (depicted by a thick green line), and $T_{p}$ (illustrated with a thin red line) are shown as functions of $h$, under fixed parameters $\mu=1.58$, $U=1$, and varying $V$ values: $0.54$ (inner curve), $0.56$ (middle curve), and $0.58$ (outer curve). For the inner curve, $T_{1}$, $T_{p^{*}}$, and $T_{p}$ are visually almost overlapping. However, as we observe the middle curve, the discrepancy between $T_{1}\approx T_{p^{*}}$ and $T_{p}$ becomes more pronounced at higher temperatures. This discrepancy is even more significant for the outer curve. Based on this discrepancy, we can conclude that the pseudo-critical transition is marked for temperatures below $T_{p}\lesssim 0.06$, where all curves overlap (when $T_{1}\approx T_{p^{*}}$ and $T_{p}$ start to diverge). A practical approach to observe the pseudo-critical temperature at higher temperatures is to consider when $\Delta T_{p}=T_{p}-T_{p^{*}}\rightarrow 0$. This is because when $\Delta T_{p}$ becomes significant, it becomes less meaningful to refer to it as a pseudo-critical temperature.
Table 2: Pseudo-critical temperature for $\mu=1.58$, $U=1$, and $V=0.56$. The second column is computed using equation (30), the third column utilizes equation (31), and the fourth column is determined by employing equation (33).
@@ IV Further Physical Quantities
Overall, physical quantities such as magnetic susceptibility or specific heat play critical roles in thermodynamics, which are essential for understanding the behavior of the Hubbard model. Therefore, we explore these physical quantities, focusing particularly on the anomalous properties they exhibit.
Figure 8: (a) Internal energy plotted against $T/T_{p}$ for various $T_{p}$ values and magnetic fields as specified in Table 2, assuming $U=1$ and $V=0.56$. (b) Entropy shown as a function of $T/T_{p}$. (c) Magnetization represented as a function of $T/T_{p}$. (d) Electron density depicted as a function of $T/T_{p}$.
To provide a clearer depiction of this anomaly, we divide the temperature $T$ by the pseudo-critical temperature $T_{p}$ listed in Table 2, where the anomaly manifests. Thus in Fig. 8a, the internal energy $\mathcal{E}$ is presented as a function of temperature normalized by the pseudo-critical temperature ($T/T_{p}$) for various fixed magnetic fields, as described within the panel. For each fixed magnetic field, a corresponding pseudo-critical temperature $T_{p}$ is given in Table 2. At a magnetic field of $h=0.0599$, the internal energy $\mathcal{E}$ remains almost constant, but there is a noticeable “jump” at $T/T_{p}=1$. This indicates a change from the quasi-alternating pair $qAP$ region to a one-electron per site region (almost polarized region) or simply a quasi-paramagnetic $qPM$ region. As the external magnetic field decreases, this sharp boundary becomes more gradual, although the “jump” in internal energy becomes larger. Similarly, in panel (b), the entropy is observed under the same conditions as in (a), highlighting these features. Notably, for a magnetic field $h\lesssim 0.06$, the boundary between $qAP$ and $qPM$ regions becomes more distinct. As $h$ approaches $0.06$, this characteristic resembles a first-order or discontinuous phase transition, despite the absence of actual discontinuity. Panel (c) depicts the magnetization $M$ as a function of $T/T_{p}$. This allows for the corroboration of the spin orientation in the $qAP$ region, where the magnetization is nearly null, while in the $qPM$ region, the magnetization is almost saturated. This feature is more pronounced at weaker magnetic fields, becoming smoother as the magnetic field decreases. Finally, panel (d) illustrates the electron density $\rho$ as a function of $T/T_{p}$, using the same set of magnetic fields considered in panel (a). Here, it is observed that the electron density remains almost constant $\rho\approx 1$ (but always smaller than $1$, $\rho<1$) for temperatures $T/T_{p}<1$, corresponding to a half-filled band. At $T/T_{p}=1$, there is a small but distinct depression in electron density, which is sharper and more pronounced at higher magnetic fields. As expected, for lower magnetic fields, the electron density decreases with increasing temperature.
@@
In Fig. 9a, the correlation length $\xi$ is illustrated as a function of $T/T_{p}$ for a fixed external magnetic field, as specified inside the panel. It is important to note that for each fixed magnetic field, there is a corresponding pseudo-critical temperature $T_{p}$, as listed in Table 2. The correlation length, $\xi$ (defined in Sec. IIB, Eq. (19), and represented by solid lines), demonstrates a decreasing function with an inflection point around $T/T_{p}=1$. As the magnetic field decreases, the inflection curvature vanishes. Contrary to what has been previously reported in the literature [1, 14, 16, 42, 41], $\xi$ does not show the typical peak observed around the pseudo-transition [1, 14, 16, 42, 41]. However, the function $1/\ln(\lambda_{0}/\lambda_{1})$ (depicted by dashed lines) exhibits a sharp peak as it approaches $h=0.06$ at $T/T_{p}=1$. The peak of this function mimics a typical peak of $\xi$ around the pseudo-transition found in references [1, 14, 16, 42, 41], indicative of a swift change between $qAP$ and $qPM$ regions. However, as the magnetic field is decreased, this sharp peak becomes more gradual. Similarly, panel (b) also illustrates the specific heat $C$ as a function of $T/T_{p}$, using the same set of magnetic fields. Here, we can observe a sharp peak, akin to that seen in a second-order phase transition, although there is no actual divergence. Furthermore, panel (c) displays the magnetic susceptibility plotted against $T/T_{p}$, clearly showing the sharp peak at $T/T_{p}=1$, which mimics a second-order phase transition. Finally, panel (d) demonstrates the behavior of isothermal compressibility $\kappa$. As $h$ approaches $0.06$, $\kappa$ diminishes, indicating that the system becomes less compressible under these conditions. Conversely, for smaller magnetic fields, $\kappa$ increases, suggesting that the system is more easily compressible. At relatively low magnetic fields, $\kappa$ exhibits a minor peak with a maximum at $T/T_{p}=1$, though this peak becomes less pronounced with lower magnetic fields.

### v2

@@ I Introduction
Recent studies on various effective-one-dimensional models with short-range interactions have revealed intriguing thermal behaviors, resembling first and second-order phase transitions [1]. This peculiar behavior to be called further pseudo-transition [1] is also dubbed ultranarrow phase crossover [2, 3], thermal pseudo-transition [4], or curious thermodynamics that resembles a phase transition [5]. Pseudo-transitions have been further analyzed in reference [6], focusing on spin correlation functions. The simplest models where this unusual behavior arises are in the Ising diamond chain [7, 3, 8]. Another similar decorated Ising chain discussed is the Ising sawtooth-like chain model [8], along with two and three-leg Ising ladder models [9, 2, 10]. Pseudo-transitions also appear in other models, such as the Ising-Heisenberg diamond chain [11, 12], and the one-dimensional double-tetrahedral model with alternating Ising spins and delocalized electrons [13]. Similar phenomena are observed in ladder models with Ising-Heisenberg coupling [14] and triangular tube models [15], highlighting a pattern of pseudo-transition. Further investigations [9, 16, 17] have shown some kind of universality of power-law exponents, while still satisfying the Rushbrooke inequality. All the above models are related to classical and classical-quantum spin models; here by classical we mean that the Hamiltonian does not contain non-commuting terms. However, models of other natures, like the extended Hubbard model in a diamond chain structure [18], and the Potts model on a diamond chain structure, have also exhibited this unusual behavior. On the other hand, exploring pseudo-transitions in genuine one-dimensional systems without decoration couplings is rather interesting. In this sense, the Zimm-Bragg-Potts model was recently explored and found to exhibit this anomalous behavior [19].
@@
On the other hand, the Hubbard model [20] stands as a foundational model in the modern theory of strongly correlated electrons, playing a paradigmatic role in the study of electronic correlations in quantum materials. This model is particularly significant in contexts where interactions are crucial. The Bethe ansatz method helps to understand this model better, especially in figuring out how electrons behave in different conditions and how they interact with each other [21]. The one-dimensional extended Hubbard model is a simplified theoretical model that describes the behavior of electrons in a one-dimensional chain, which additionally includes nearest-neighbor interaction energy terms. The nearest-neighbor interaction energy term describes the Coulomb repulsion between electrons occupying neighboring sites. The one-dimensional extended Hubbard model has been widely studied in the literature due to its relevance in understanding the electronic properties of one-dimensional materials such as carbon nanotubes and organic conductors [22]. It also serves as a simple prototype model for studying strongly correlated electron systems [23]. On the other hand, several theoretical investigations have focused on the extended Hubbard model in the one-dimensional case. Numerous investigations regarding the half-filled ground states of the extended Hubbard model have been conducted. Tsuchiizu and Furusaki [24] revisited the ground-state phase diagram of the one-dimensional half-filled extended Hubbard model, revealing different phases and transitions using a renormalization-group approach. Glocke et al. [25] utilized density-matrix renormalization group methods on transfer matrices to study the thermodynamics of the one-dimensional extended Hubbard model at half-filling. This highlights the detection of phase transitions through standard thermodynamic measures like isothermal compressibility and magnetic susceptibility. The study identifies a unique phase with long-range dimer order and delineates the phase diagram, comparing it with quantum Monte Carlo studies [26, 27]. The phase diagram and power-law exponents of the one-dimensional $U-V$ model at quarter-filling were identified using exact diagonalization and various limit results, identifying a transition from Luttinger liquid to charge density wave insulator, noting dominant superconducting or spin density wave fluctuations depending on the value of $V$ [28]. Further investigation was also performed in reference [29].
@@
The present work is organized as follows: In Sec. 2, we give the thermodynamics of the one-dimensional extended Hubbard model in the atomic limit, identifying each eigenvalue of the cubic root and determining which one is the largest, a topic not previously elucidated. In Sec. 3, we analyze a peculiar property in the low-temperature region and explore the anomalous behavior, along with the corresponding region where the main properties undergo significant changes at this anomalous temperature. This phenomenon is what we refer to as the quasi-phase diagram [1]. In Sec. 4, we explore additional physical quantities, reporting the influence of pseudo-transition. Finally, in Sec. 5, we present our conclusions.
@@ II.2 Thermodynamic quantities and correlators
Here, we introduce several useful thermodynamic quantities; all these quantities will be quantified per site. The entropy is calculated as ${\cal S}=-\frac{\partial\Omega}{\partial T}$; the enthalpy is given by $\mathcal{E}=-\frac{\partial\ln\left(\lambda_{0}\right)}{\partial\beta}=k_{B}T^{2}\frac{\partial\ln\left(\lambda_{0}\right)}{\partial T}$; analogously we can obtain specific heat at constant chemical potential $C=T\frac{\partial{\cal S}}{\partial T}=\frac{\partial{\cal E}}{\partial T}$ [40]; the magnetization can be expressed as $M=-\frac{\partial\Omega}{\partial h}$; magnetic susceptibility $\chi=\frac{\partial M}{\partial h}$; the electron density $\rho$ is determined by $\rho=-\frac{\partial\Omega}{\partial\mu}$; and the isothermal compressibility is derived from $\kappa=\tfrac{1}{\rho^{2}}\frac{\partial\rho}{\partial\mu}$.
@@ III Phase diagrams
In what follows, we discuss the ground-state energy, focusing on peculiar regions where anomalous behavior appears. Then we consider the low-temperature case to discuss quasi-phase diagrams and define a peculiar temperature.
@@ III.1 Zero-temperature phase diagram
Table 1: The second column provides zero-temperature phase boundary conditions, the third column gives the chemical potential at the phase boundary, while the fourth column reports the associated residual boundary entropy. Note that for the $FR_{2}-AP$ there is no dependence on $\mu$ according to the second column and therefore the corresponding row in the third column is empty.
@@ III.2 Anomalous behavior in low-temperature region and quasi-phase diagram
Figure 5: Quasi-phase diagram depicting the half-filled band in the $h/U-T/U$ plane, with fixed parameters $V/U=0.54$ and $\mu/U=1.58$. The dashed curve delineates the peculiar temperature $T_{p}$. Panels (a) and (b) showcase density plots of magnetization and entropy, respectively.
@@
In exploring another intriguing aspect of the low-temperature region, Fig. 5a presents a density plot of magnetization $M$ in the $h/U-T/U$ plane, assuming $V/U=0.54$ and $\mu/U=1.58$. Here, the dashed line represents the peculiar temperature $T_{p}$. Below a magnetic field of $h/U=0.04$ at sufficiently low temperatures, the system resides in a quasi alternating-pair $qAP$ region, characterized by particles predominantly arranged in alternating pairs. As the magnetic field increases within this low temperature regime, the system goes swiftly into a quasi-paramagnetic $qPM$ region, where most spins, one per site, align parallel to the magnetic field. Conversely, at very weak magnetic fields and increasing temperatures, the $FR_{2}$ frustrated system leads into a quasi-frustrated $qFR_{2}$ region. The boundary between $qFR_{2}$ and $qPM$ is marked by a standard crossover, or as visualized in the magnetization plot, a diffuse boundary, appearing as a fuzzy region. In panel (a), one can observe the magnetization shift between $qAP$ and $qPM$ regions, although the boundary between $qAP$ and $qFR_{2}$ remains indistinct due to null magnetization in both regions. Similarly, Fig. 5b displays the entropy $\mathcal{S}$, using the same parameters as panel (a). Here, the crossover boundary, denoted by a dashed line, outlines the peculiar temperature as a function of magnetic field. In contrast, panel (b) shows that the boundary between $qAP$ and $qPM$ becomes indistinguishable with almost null entropy, while the boundary between $qAP$ and $qFR_{2}$ is sharply defined. This distinction is due to the residual entropy of $FR_{2}$ being $\mathcal{S}=k_{B}\ln(2)$. However, at temperatures above $T>0.06$, the entropy significantly increases, becoming clearly larger than $\mathcal{S}\gtrsim 0.7k_{B}$.
III.3 Peculiar temperature condition
@@
This detailed analysis further explores the interesting behavior in the low-temperature region of the model. Previously, the peculiar temperature was defined for models where the free energy involved a simple square root expression [1, 6, 16]. Defining the peculiar temperature generally poses a challenge, as the only available indicator is an anomalous behavior in the low-temperature region, marked by significant changes in entropy and magnetization, along with sharp peaks in correlation length, specific heat, and magnetic susceptibility. At very low temperatures, these peaks occur at roughly the same temperature. However, as the anomalous behavior shifts to higher temperatures $T_{p}$, these peaks appear at slightly different temperatures for each quantity, complicating the definition of the peculiar temperature.
@@
An interesting observation is that this anomalous behavior is evident in fundamental quantities such as the ratio between the second largest and the largest eigenvalues. To analyze this more deeply, consider the temperature and magnetic field dependence of the eigenvalues. At a fixed magnetic field $h_{1}$, there exists a special temperature, the peculiar temperature $T_{p}$, which satisfies the relation
@@
Panel (b) of the Fig. 6 illustrates the function $g(T,h)$ as it varies with temperature, under the conditions of a null magnetic field, $\mu=1.58$, $U=1$, and $V=0.55$. In this representation, one can observe that the function $g(T,h)$ passes through zero at a specific temperature, denoted as $T_{1}$. This provides a clear visual indication of the peculiar point in the behavior of the system.
@@
Furthermore, by employing a perturbative approach, as elaborated in Appendix A and illustrated through Eqs. (41) and (42) for the unperturbed result, we reaffirm the condition of the equation derived earlier, simplifying our findings into this coherent expression.
In regions of low temperature, $T_{p}\approx T_{p^{*}}\approx T_{1}$, which are very close to each other. When this occurs, we can observe anomalous behavior, and we simply define it as the peculiar temperature, $T_{p}$.
For the present model the condition (36) provides a sufficiently accurate condition for determining the peculiar temperature, where $T_{p}\approx T_{1}$ holds. Specifically, the peculiar temperature for a fixed magnetic field $h_{1}$ can be represented using the results $w_{0,2}={\rm e}^{\beta(\mu-U/2)}$, $w_{1,1}={\rm e}^{\beta(\mu-V)}$, and $z=2\cosh(\beta h)$, leading to the condition
@@
In Fig. 7, the temperatures $T_{1}$ (represented by a thin black line), $T_{p^{*}}$ (depicted by a thick green line), and $T_{p}$ (illustrated with a thin red line) are shown as functions of $h$, under fixed parameters $\mu=1.58$, $U=1$, and varying $V$ values: $0.54$ (inner curve), $0.56$ (middle curve), and $0.58$ (outer curve). For the inner curve, $T_{1}$, $T_{p^{*}}$, and $T_{p}$ are visually almost overlapping. However, as we observe the middle curve, the discrepancy between $T_{1}\approx T_{p^{*}}$ and $T_{p}$ becomes more pronounced at higher temperatures. This discrepancy is even more significant for the outer curve. Based on this discrepancy, we can conclude that the peculiar transition is marked for temperatures below $T_{p}\lesssim 0.06$, where all curves overlap (when $T_{1}\approx T_{p^{*}}$ and $T_{p}$ start to diverge). A practical approach to observe the peculiar temperature at higher temperatures is to consider when $\Delta T_{p}=T_{p}-T_{p^{*}}\rightarrow 0$. This is because when $\Delta T_{p}$ becomes significant, it becomes less meaningful to refer to it as a peculiar temperature.
Table 2: Peculiar temperature for $\mu=1.58$, $U=1$, and $V=0.56$. The second column is computed using equation (30), the third column utilizes equation (31), and the fourth column is determined by employing equation (33).
@@ IV Further Physical Quantities
Overall, physical quantities such as magnetic susceptibility or specific heat play fundamental roles in thermodynamics, which are essential for understanding the behavior of the Hubbard model. Therefore, we explore these physical quantities, focusing particularly on the anomalous properties they exhibit.
Figure 8: (a) Enthalpy plotted against $T/T_{p}$ for various $T_{p}$ values and magnetic fields as specified in Table 2, assuming $U=1$ and $V=0.56$. (b) Entropy shown as a function of $T/T_{p}$. (c) Magnetization represented as a function of $T/T_{p}$. (d) Electron density depicted as a function of $T/T_{p}$.
To provide a clearer depiction of this anomaly, we divide the temperature $T$ by the peculiar temperature $T_{p}$ listed in Table 2, where the anomaly manifests. Thus in Fig. 8a, the enthalpy $\mathcal{E}$ is presented as a function of temperature normalized by the peculiar temperature ($T/T_{p}$) for various fixed magnetic fields, as described within the panel. For each fixed magnetic field, a corresponding peculiar temperature $T_{p}$ is given in Table 2. At a magnetic field of $h=0.0599$, the enthalpy $\mathcal{E}$ remains almost constant, but there is a noticeable “jump” at $T/T_{p}=1$. This indicates a change from the quasi-alternating pair $qAP$ region to a one-electron per site region (almost polarized region) or simply a quasi-paramagnetic $qPM$ region. As the external magnetic field decreases, this sharp boundary becomes more gradual, although the “jump” in enthalpy becomes larger. Similarly, in panel (b), the entropy is observed under the same conditions as in (a), highlighting these features. Notably, for a magnetic field $h\lesssim 0.06$, the boundary between $qAP$ and $qPM$ regions becomes more distinct. As $h$ approaches $0.06$, this characteristic resembles a first-order or discontinuous phase transition, despite the absence of actual discontinuity. Panel (c) depicts the magnetization $M$ as a function of $T/T_{p}$. This allows for the corroboration of the spin orientation in the $qAP$ region, where the magnetization is nearly null, while in the $qPM$ region, the magnetization is almost saturated. This feature is more pronounced at weaker magnetic fields, becoming smoother as the magnetic field decreases. Finally, panel (d) illustrates the electron density $\rho$ as a function of $T/T_{p}$, using the same set of magnetic fields considered in panel (a). Here, it is observed that the electron density remains almost constant $\rho\approx 1$ (but always smaller than $1$, $\rho<1$) for temperatures $T/T_{p}<1$, corresponding to a half-filled band. At $T/T_{p}=1$, there is a small but distinct depression in electron density, which is sharper and more pronounced at higher magnetic fields. As expected, for lower magnetic fields, the electron density decreases with increasing temperature.
@@
In Fig. 9a, the correlation length $\xi$ is illustrated as a function of $T/T_{p}$ for a fixed external magnetic field, as specified inside the panel. It is important to note that for each fixed magnetic field, there is a corresponding peculiar temperature $T_{p}$, as listed in Table 2. The correlation length, $\xi$ (defined in Sec. IIB, Eq. (19), and represented by solid lines), demonstrates a decreasing function with an inflection point around $T/T_{p}=1$. As the magnetic field decreases, the inflection curvature vanishes. Contrary to what has been previously reported in the literature [1, 16, 18, 44, 43], $\xi$ does not show the typical peak observed around the pseudo-transition [1, 16, 18, 44, 43]. However, the function $1/\ln(\lambda_{0}/\lambda_{1})$ (depicted by dashed lines) exhibits a sharp peak as it approaches $h=0.06$ at $T/T_{p}=1$. The peak of this function mimics a typical peak of $\xi$ around the pseudo-transition found in references [1, 16, 18, 44, 43], indicative of a swift change between $qAP$ and $qPM$ regions. However, as the magnetic field is decreased, this sharp peak becomes more gradual. Similarly, panel (b) also illustrates the specific heat $C$ as a function of $T/T_{p}$, using the same set of magnetic fields. Here, we can observe a sharp peak, akin to that seen in a second-order phase transition, although there is no actual divergence. Furthermore, panel (c) displays the magnetic susceptibility plotted against $T/T_{p}$, clearly showing the sharp peak at $T/T_{p}=1$, which mimics a second-order phase transition. Finally, panel (d) demonstrates the behavior of isothermal compressibility $\kappa$. As $h$ approaches $0.06$, $\kappa$ diminishes, indicating that the system becomes less compressible under these conditions. Conversely, for smaller magnetic fields, $\kappa$ increases, suggesting that the system is more easily compressible. At relatively low magnetic fields, $\kappa$ exhibits a minor peak with a maximum at $T/T_{p}=1$, though this peak becomes less pronounced with lower magnetic fields.